OpenStudy (anonymous):

I need help answering these questions for Platinum I need to find the moles of each element and atoms of each element for platinum.

OpenStudy (abb0t):

Given a molar number of some element, you are going to use dimensional analysis to find the moles and or grams. That means you are going to be using your periodic table for this. For example, say in an experiment, you have 2 grams of Pt (platinum). To find the moles, you are basically cross canceling out like this (since you know that the units for the elements in the periodic table are in units of: grams per mol). \[2 g. Pt \times \frac{ 1 mol }{ 195 g. Pt } = 2 \div 195 mol. Pt\]notice how the two units of grams cancel out. It's similar to algebra where you may have: \[\frac{ a }{ x } \times \frac{ b }{ a } = \frac{ b }{ x }\] that two a's cancel out since you know that anything that is \[\frac{ a }{ a } = 1\] Then to convert or find atoms, you are going to use avogadro's number, which I think is \[6.022 \times 10^{23}\]. So you are going to use dimensional analysis again. The format is: \[\frac{ 2\div 195 }{ 6.022 \times 10^{23} } = atoms. of. Pt\]